I'm not sure what type of problem this is :( Please someone help me.
The ratio of two #'s is 8 to 5. Also, 3 times the larger # is 3 less that 5 times the smaller #. What are the #'s?

Thanks.

Answer by askmemath(368)

Let the 2 numbers be 8x and 5x
3 times the larger = 3(8x) = 24x
is 3 less than 5 times the smaller
24X = 5(5x)-3
24x = 25x-3
Simplifying X = 3
So the numbers are 24 and 15
Check:
3 times the larger = 3(24) = 72
5 times the smaller = 5(15) = 75
Difference = 75-72 = 3
Hence the answer is correct :-)